OpenStudy (anonymous):

A building lot in a city is shaped as a 30-60-90 triangle. The side opposite the 30 angle measures 41 feet. a. Find the length of the side of the lot opposite the 60 angle. b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30 angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (anonymous):

I've got B I just need help on A and C

OpenStudy (anonymous):

ratios of a 30-60-90 right triangle are \(1:\sqrt{3}:2\) for the short side, long side and hypotenuse respectively

OpenStudy (anonymous):



OpenStudy (anonymous):

so short side is 41, long side is \(41\times \sqrt{3}\) and hypotenuse is \(2\times 41=82\)

OpenStudy (anonymous):

Alright, I think I've got it now! Thank you so much!

OpenStudy (anonymous):

sine , cosine and tangent are not dependent on the lengths of the sides of the triangle, only the rations \(\sin(30)=\frac{1}{2}, \cos(30)=\frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

*ratios not "rations" yw